5. A balloon operates at an altitude where the air density is 0.9 kg / m3. At that altitude, the balloon has a volume of 20m3; the balloon is filled with helium (helium has a density of: ρHe = 0178 kg / m3). If the weight of the balloon bag is 94N, what load can the globe withstand at this altitude?
1
Expert's answer
2018-09-08T16:55:08-0400
For the equilibrium:
W+F=(ρ_air-ρ_He )gV
F=(ρ_air-ρ_He )gV-W
F=(0.9-0.178)(9.8)(20)-94= 47.5 N.